OpenStudy (anonymous):

If 7p + 5q = 2 and 8p - 9q = 17, then -p + 14q =

OpenStudy (anonymous):

set up a system of equations with the first two equations. 7p + 5q = 2 8p -9q = 17 --------------- solve for either p or q. then solve for the other variable. use the two values to solve the equation (-p+14q)

OpenStudy (anonymous):

how do i solve for p or q?

OpenStudy (anonymous):

7p + 5q = 2 8p -9q = 17 --------------- multiply the top by 9 and the bottom by 5 63p + 45q = 18 40p - 45q = 85 ----------------- 103p + 0q = 103 103p = 103 p = 1 use equation (7p+5q = 2) to solve for q 7(1) + 5q = 2 5q = 2-7 5q = -5 q = -1 try solving the final equation with these values.